2 coin lie 2.3m apart in vacuum. they carry identical charges. how long is the charge on each if a coin experience a force of 4.5N.(q=51.4C)
1
Expert's answer
2021-08-20T12:55:46-0400
"F=k\frac{q^2}{r^2}\\4.5=(9\cdot10^9)\frac{q^2}{2.3^2}\\q=51.4\ \mu C"